-15 apart. 1: Calculate the electric force between 2 protons in a nucleus that are 5x101 m Compare this with the gravitational force between them. The gravitational force is expressed by F-GMr where M, and M2 are the masses of the two objects and G is the gravitational constant 6.67x101Nm2kg?. Your answers here might make you wonder why the protons dont fly apart. Use m,-1.67x102/kg. -27
Answer #1

Electric force is given by:

Fe = k*q1*q2/r^2

q1 = q2 = charge on proton = +e = 1.6*10^-19 C

k = electrostatic constant = 9*10^9

r = distance between protons = 5*10^-15 m

So,

Fe = 9*10^9*(1.6*10^-19)^2/(5*10^-15)^2

Fe = 9.22 N

Now Gravitational force is given by:

Fg = G*m1*m2/r^2

m1 = m2 = mass of proton = 1.67*10^-27 kg

So,

Fg = 6.67*10^-11*(1.67*10^-27)^2/(5*10^-15)^2

Fg = 7.44*10^-36 N

Comparison between Fe and Fg

Fe/Fg = 9.22/(7.44*10^-36)

Fe = 1.24*10^36*Fg

Repulsive Electric force between proton is 10^36 times the attractive gravitational force between protons.

Please Upvote.